40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6321     Entry time: Sat Feb 25 14:27:26 2012
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: glitches in the RFPD outputs 

Last night I took a closer look at the LSC analog signals to find which components are making the glitches.

I monitored the RFPD output signals and the demodulated signals at the same time with an oscilloscope when the PRMI was kept locked.

Indeed the RFPD outputs have some corresponding fast signals although I only looked at the RELL11 I and Q signals.

(REFL33 didn't have sufficiently a high SNR to see the glitches with the oscilloscope.)

I will check the rest of channels.

ELOG V3.1.3-